OpenStudy (anonymous):

Fan And Medal! Factor Completely: x2 + 25 (x + 5)(x + 5) (x + 5)(x − 5) (x − 5)(x − 5) Prime

OpenStudy (anonymous):

(x + 5)(x + 5)

OpenStudy (anonymous):

@SkydexKiwi that's wrong

OpenStudy (anonymous):

You are missing one term. This expression is not factorable

OpenStudy (anonymous):

what!?

OpenStudy (anonymous):

You need to use complex solutions.

OpenStudy (anonymous):

So its a prime?

OpenStudy (anonymous):

@robz8 please explain?

OpenStudy (anonymous):

oh wait it' (sx + 5)(x − 5) isn't it?

OpenStudy (psymon):

How is it not prime?

OpenStudy (anonymous):

(x + 5)(x − 5)

OpenStudy (anonymous):

it is prime

OpenStudy (psymon):

@SkydexKiwi Thats the factored form of x^2 - 25

OpenStudy (anonymous):

haha i suck then been a while for this

OpenStudy (anonymous):

You could do (a +b)^2 -2ab)

OpenStudy (anonymous):



OpenStudy (mathstudent55):

The difference of squares is factorable. The sum of squares is not factorable. The difference of cubes is factorabnle. The sum of cubes is factorable. This is a sum of squares, so it's not factorable.

OpenStudy (anonymous):

x2 + 25? If x = 2 for example then x2 + 25 = 4 + 25 = 29 which is a prime number. Thus the only factors of 29 are 29 and 1 and hence the only factorization of 29 is 29 × 1. Similarly if x = 4 then x2 + 25 = 41 which is again prime. Hence the only factorization of x2 + 25 which is valid for every number x is

OpenStudy (anonymous):

thats wha t i got from another site so i think im going with robs

OpenStudy (anonymous):

You could also do (a+bi)(a-bi)... which is not a valid factorization... I knew it!! @robz8 ...well ?

OpenStudy (anonymous):